What is the most we should pay for a bond with a par value of $1000, coupon rate of 7.3% paid annually, and a remaining life of 26 years? The yield to maturity is 10.0%. Assume annual discounting. (Round your answer to the nearest penny.)
Answer #1

Annual coupon=1000*7.3%=73

Hence current price=Annual coupon*Present value of annuity factor(10%,26)+$1000*Present value of discounting factor(10%,26)

=73*9.16094547+1000*0.0839054529

=$752.65(Approx).

NOTE:

1.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=73[1-(1.1)^-26]/0.1

=73*9.16094547

2.Present value of discounting factor=1000/1.1^26

=1000*0.0839054529